OpenStudy (narissa):

a flag pole casts a shadow that is 30 feet in length. A roope that is 40 feet in length is tied to the top and is extended to the edge of the shadow. Find the height of the flag pole to the nearest tenth.

OpenStudy (narissa):

@umerlodhi

OpenStudy (narissa):

do u know pythagorean theorem

OpenStudy (narissa):

it says x=26.5 ft but i have to show how to get that

umerlodhi (umerlodhi):

yes i learned it last year

umerlodhi (umerlodhi):

Basically it is a^2 + b^@=c^2 30*30=900 40*40=1600 2500 √2500=50

umerlodhi (umerlodhi):

new thread pls